OpenStudy (anonymous):

Last question! When you factor completely, the factors of 16-x^4 are: ?

OpenStudy (anonymous):

difference of 2 squares: (4-x^2)(4+x^2) (2-x)(2+x)(4+x^2)

OpenStudy (anonymous):

do you want imaginary number factors as well?

OpenStudy (anonymous):

if you include imaginary numbers, fully factored would be: (2-x)(2+x)(x-2i)(x+2i)

OpenStudy (anonymous):

(4-x^2)(4+x^2) are the solutions so, 4-x^2=0 4=x^2 + 2 or -2=x 4+x^2=0 x^2=-4 x= square root -4 or 2i

OpenStudy (anonymous):

You guys are awesome. Thanks!